OpenStudy (anonymous):

express in index form: (2n)^4÷8n^0

OpenStudy (sepeario):

ok lets firstly get things straight.

OpenStudy (sepeario):

sorry \[8n ^{0}=8*1\]

OpenStudy (sepeario):

because n^0 = 1

OpenStudy (sepeario):

hence the equation is \[(2n^4)/8 = 2n^4/2^3\]

OpenStudy (sepeario):

BUT

OpenStudy (sepeario):

there is still one more step, divide both top and bottom by 2

OpenStudy (sepeario):

\[n^4/2^{2} = n^4/4\]

OpenStudy (sepeario):

sorry about the long explanation lol

OpenStudy (anonymous):

u have a mistake from the beginning: it should be 16n^4/8=2n^4

OpenStudy (anonymous):

wow. thanks